If a triangle has a base of 11’ and a height of 6’, what is the length of the hypotenuse?


Answer by checkley79(3341)

IF this is a right triangle then:
6^6+11^2=c^2
36+121=c^2
c^2=157
c=sqrt157
c=12.5' ans. for the hypotenuse.
IF this is not a right triangle then the answer is between 12.32' & 7.78'.
5.5^2+6^2=c^2
30.25+36=c^2
c^2=66.25
c=sqrt66.25
c=7.78' ans.